### v2

@@ Abstract
We consider a dynamic model of traffic that has received a lot of attention in the past few years. Users control infinitesimal flow particles aiming to travel from an origin to a destination as quickly as possible. Flow patterns vary over time, and congestion effects are modeled via queues, which form whenever the inflow into a link exceeds its capacity. Despite lots of interest, some very basic questions remain open in this model. We resolve a number of them:
@@ 1 Introduction
Consider a network of such links, and suppose that all traffic in the network has the same origin and destination. Clearly this is a restrictive assumption, but already this restricted setting is very challenging (moreover, we will argue in Section 7 that our results extend to the more general single-commodity setting where multiple sources and sinks are allowed). Starting at time 00, users are released into the system at the origin at some constant rate; each user controls an infinitesimal amount of flow. Their joint choices yield a flow over time, as first introduced by Ford and Fulkerson in the fifties [FF56]. But here, each user is self-interested and wants to find a quickest route to the destination. This will depend on the choices of the other users, since these will impact the queue lengths. A Nash equilibrium in this model is then a joint choice of routes for all users, such that all users are satisfied in hindsight with their choices: no user can switch to an alternative route and arrive at the destination at a strictly earlier time. Note that this means that when making a decision, it is not the queueing delays now that matter, but the queues that the user would actually experience upon reaching an arc, which may be different. An alternative notion of “instantaneous dynamic equilibria” [GHS20] has also been considered, where users base decisions on current queue lengths rather than what they will actually experience, and more recent work by Graf, Harks, Kollias and Merkl [GHKM22] considers a much more general information framework where users use predictions of future congestion patterns. This paper concerns only Nash equilibria.
@@
There has been some substantial progress in understanding equilibria in this model. Koch and Skutella [KS11] (see also [Koc12]) were the first to study it using tools from combinatorial optimization. They uncovered an intricate structure in the time derivatives of equilibria, which they could precisely describe as a solution to a certain nonlinear system (which they called thin flow with resetting, and which we will encounter later). Assuming solutions to this system always exist, they showed how an equilibrium could be constructed by essentially integrating; the resulting equilibrium has the property that all queue lengths are piecewise linear. They call a maximal interval of departure times on which queues are affine a phase. (Note that there is no guarantee that there are a finite number of phases, even within a compact interval; that is, the algorithm is not known to be finite).
@@
Existence was shown with precision by Cominetti, Correa and Larré [CCL15]. They showed that the nonlinear system of Koch and Skutella always has a solution (in fact a unique solution), and hence that the integration procedure always succeeds in constructing an equilibrium. Surprisingly, it is an open question as to whether the thin flow equations can be efficiently solved; in other words, whether there is a polynomial-time algorithm to compute the next phase, given that the behavior for all previous phases has been computed.
@@
[CCL15] give a partial answer to this question. They show that if one restricts to “right differentiable” solutions (the precise meaning of this we postpone until later), then equilibria are unique with respect to journey times. First impressions might be that such a restriction is of a technical nature, and that this essentially shows uniqueness, but this is misleading. The precise reason for this requires more technical preliminaries, and we postpone this until Section 3. For now, we remark that there is no a priori reason why equilibria not respecting this right-differentiability requirement could not exist, nor anything “unphysical” about them.
We prove that this cannot happen, and that equilibria are indeed unique, without any assumptions. To further motivate why this uniqueness result is important, we will shortly discuss how it relates to the next question we resolve.
@@ 2 Model and preliminaries
An instance is described by a directed graph $G=(V,E)$ with a source $s$ and sink $t$, where each arc is equipped with a transit time $\tau_{e}\geq 0$ and a capacity $\nu_{e}>0$. At $s$ we have a constant network inflow rate of $u_{0}>0$, which begins at time 00. We may assume that every node in $G$ is both reachable from $s$, and can reach $t$. For technical convenience, we will follow previous works and assume that $G$ has no directed cycle consisting of arcs with transit time 00.
We use the notation $\delta^{-}(v)$ and $\delta^{+}(v)$ to denote the set of incoming and outgoing arcs at $v$, respectively, and similarly $\delta^{-}(S)$ and $\delta^{+}(S)$ for arcs entering or leaving a set $S$. We define $[z]^{+}\coloneqq\max\{0,z\}$.
@@
The earliest arrival time at $w$ of a particle starting at $s$ at time $\theta\in\mathbb{R}_{\geq 0}$ is given by
@@
We will often refer to the collection of earliest arrival times for a fixed $\theta$ as a labeling. Given earliest arrival times $\ell$, we define the associated queueing delays (or queueing delay functions) by $q_{e}(\theta)\coloneqq z_{e}(\ell_{v}(\theta))/\nu_{e}$ for each $e=vw\in E$. Note that $q_{e}(\theta)$ is the queueing delay on arc $e$ for a flow particle departing $s$ at time $\theta$ and taking a shortest path to $e$. For any earliest arrival times $\ell$, we define the active arcs $E^{\prime}_{\ell(\theta)}$ and the resetting arcs $E^{*}_{\ell(\theta)}$ for some time $\theta$ as
$E^{\prime}_{\ell(\theta)}\coloneqq\set{e=vw\in E}{\ell_{w}(\theta)=\ell_{v}(\theta)+q_{e}(\theta)+\tau_{e}}\quad\text{and}\quad E^{*}_{\ell(\theta)}\coloneqq\set{e=vw\in E}{q_{e}(\theta)>0}$
@@
An equilibrium (also referred to as a dynamic equilibrium or a Nash flow over time) is a feasible flow over time in which almost all flow particles travel along a shortest path from $s$ to $t$, i.e., only along active arcs. In this case the active and resetting arcs are characterized by
$E^{\prime}_{\ell(\theta)}=\set{e=vw\in E}{\ell_{w}(\theta)\geq\ell_{v}(\theta)+\tau_{e}}\quad\text{ and }\quad E^{*}_{\ell(\theta)}=\set{e\in E}{\ell_{w}(\theta)>\ell_{v}(\theta)+\tau_{e}}.$ (1)
@@
As proven in [CCL15], a feasible flow over time is an equilibrium if and only if $F^{+}_{e}(\ell_{v}(\theta))=F^{-}_{e}(\ell_{w}(\theta))$ for all arcs $e=vw$ and all departure times $\theta$. Define $x_{e}(\theta)\coloneqq F^{+}_{e}(\ell_{v}(\theta))$ for all $e$ and $\theta$. For an equilibrium flow, the derivative of $x_{e}$ at time $\theta$, which exists almost everywhere, can be interpreted as a flow describing what proportion of flow particles departing at time $\theta$ use arc $e$; $x^{\prime}$ is an $s$-$t$-flow of value $u_{0}$ almost everywhere. It has been shown [CCL15, KS11] that a flow over time is an equilibrium if and only if the resulting pair $(x,\ell)$ satisfies the following for almost every $\theta$: setting $x^{\prime}=x^{\prime}(\theta)$, $\ell^{\prime}=\ell^{\prime}(\theta)$, $E^{\prime}=E^{\prime}_{\theta}$ and $E^{*}=E^{*}_{\theta}$,
$x^{\prime}$ is a static $s$-$t$-flow of value $u_{0}$, (TF-1)

### v3

@@ Abstract
We consider a dynamic model of traffic that has received a lot of attention in the past few years. Users control infinitesimal flow particles aiming to travel from an origin to a destination as quickly as possible. Flow patterns vary over time, and congestion effects are modeled via queues, which form whenever the inflow into a link exceeds its capacity. Despite lots of interest, some very basic questions remain open in this model. We resolve a number of them in the single-commodity setting:
@@ 1 Introduction
Consider a network of such links, and suppose that all traffic in the network has the same origin and destination. Clearly this is a restrictive assumption, but already this restricted setting is very challenging (moreover, we will argue in Section 7 that our results extend to the more general single-commodity setting where multiple sources and sinks are allowed). Starting at time 0, users are released into the system at the origin at some constant rate; each user controls an infinitesimal amount of flow. Their joint choices yield a flow over time, as first introduced by Ford and Fulkerson in the fifties [FF56]. But here, each user is self-interested and wants to find a quickest route to the destination. This will depend on the choices of the other users, since these will impact the queue lengths. A Nash equilibrium in this model is then a joint choice of routes for all users, such that all users are satisfied in hindsight with their choices: no user can switch to an alternative route and arrive at the destination at a strictly earlier time. Note that this means that when making a decision, it is not the queueing delays now that matter, but the queues that the user would actually experience upon reaching an arc, which may be different. An alternative notion of “instantaneous dynamic equilibria” [GHS20] has also been considered, where users base decisions on current queue lengths rather than what they will actually experience, and more recent work by Graf, Harks, Kollias and Merkl [GHKM22] considers a much more general information framework where users use predictions of future congestion patterns. This paper concerns only Nash equilibria.
@@
There has been some substantial progress in understanding equilibria in this model. Koch and Skutella [KS11] (see also [Koc12]) were the first to study it using tools from combinatorial optimization. They uncovered an intricate structure in the time derivatives of equilibria, which they could precisely describe as a solution to a certain nonlinear system (which they called thin flow with resetting, and which we will encounter later). Assuming solutions to this system always exist, they showed how an equilibrium could be constructed by essentially integrating; the resulting equilibrium has the property that all queue lengths are piecewise linear. They call a maximal interval of departure times on which queues are affine a phase. (Note that there is no guarantee that there are a finite number of phases, even within a compact interval; that is, the algorithm is not known to be finite). A similar integration approach can be found in the transportation literature; see Kuwahara and Akamatsu [KA93] and the survey by Iryo [Iry13].
@@
Existence of equilibria is a crucial question. Zhu and Marcotte [ZM00] showed existence of equilibria for a somewhat general model where congestion satisfies a “strong FIFO” property; however, the fluid queing model does not have this property. Mounce [Mou06] showed existence in the fluid queueing model if each possible route a user may take contains at most one bottleneck arc of finite capacity. A general existence result, without any such restrictions, was shown by Cominetti, Correa and Larré [CCL15]. Results for models with departure time choice and more general congestion behavior were obtained by Meunier and Wagner [MW10] and Han, Friesz and Yao [HFY13]. Here we focus our discussion on the approach taken by [CCL15], as it will be most relevant to our approach. They showed that the nonlinear system of Koch and Skutella always has a solution (in fact a unique solution), and hence that the integration procedure always succeeds in constructing an equilibrium. Surprisingly, it is an open question as to whether the thin flow equations can be efficiently solved [^1] The only known positive result is an algorithm on series-parallel graphs [Kai20].; in other words, whether there is a polynomial-time algorithm to compute the next phase, given that the behavior for all previous phases has been computed.
@@
This question has been discussed extensively in the literature. While we are focused here on the setting of a single origin and destination, Iryo [Iry11] showed that it can fail to hold if there are multiple origins and destination (though Iryo and Smith give positive results for “unidirectional” networks [IS18]). Uniqueness has been demonstrated in the single-origin single-destination setting under additional assumptions of various strengths, e.g., [KA93, SM07, CCL15]. But to the best of our knowledge, none of these works fully address the question for the bottleneck model without any additional assumptions. A somewhat subtle — but as we will see, very important — aspect of this is explained clearly in [CCL15]. They give a proof of uniqueness — but only if one restricts to right-differentiable solutions. All other works based on extending an equilibrium over time, to the best of our knowledge, explicitly or implicitly require this right-differentiability assumption. First impressions might be that such a restriction is of a technical nature, and that this essentially shows uniqueness, but this is misleading. The precise reason for this requires more technical preliminaries, and we postpone this until Section 3. For now, we remark that there is no a priori reason why equilibria not respecting this right-differentiability requirement could not exist, nor anything “unphysical” about them.
We prove that equilibria are indeed unique in the single-origin single-destination setting with constant inflow, without such assumptions. To further motivate why this uniqueness result is important, we will shortly discuss how it relates to the next question we resolve.
@@
Piecewise-constant inflows.
While we have described the model and our results in the context of constant inflow, it is rather straightforward to see that continuity and uniqueness of equilibria with piecewise-constant inflows is an immediate consequence.
Of course, our result on reaching steady state does not hold for piecewise-constant inflows in general — if the inflow is varying forever, we cannot expect queues to behave linearly.
@@ 2 Model and preliminaries
An instance is described by a directed graph $G=(V,E)$ with a source $s$ and sink $t$, where each arc is equipped with a transit time $\tau_{e}\geq 0$ and a capacity $\nu_{e}>0$. At $s$ we have a constant network inflow rate of $u_{0}>0$, which begins at time 0. We may assume that every node in $G$ is both reachable from $s$, and can reach $t$. For technical convenience, we will follow previous works and assume that $G$ has no directed cycle consisting of arcs with transit time 0.
We use the notation $\delta^{-}(v)$ and $\delta^{+}(v)$ to denote the set of incoming and outgoing arcs at $v$, respectively, and similarly $\delta^{-}(S)$ and $\delta^{+}(S)$ for arcs entering or leaving a set $S$. Moreover, let $E[S]$ denote arcs with both end points in a set $S$. We define $z]^{+}\coloneqq\max\{0,z\}$.
2.1 Flows over time
@@
We will start at time 0 with an empty network, requiring that in a feasible flow $f_{e}^{-}(\theta)=0$ for all $\theta<\tau_{e}$ and all $e\in E$. The earliest arrival time at $w$ of a particle starting at $s$ at time $\theta\in\mathbb{R}_{\geq 0}$ is given by
@@
We will often refer to the collection of earliest arrival times for a fixed $\theta$ as a labeling. Given earliest arrival times $\ell$, we define the associated queueing delays (or queueing delay functions) by $q_{e}(\theta)\coloneqq z_{e}(\ell_{v}(\theta))/\nu_{e}$ for each $e=vw\in E$. Note that $q_{e}(\theta)$ is the queueing delay on arc $e$ for a flow particle departing $s$ at time $\theta$ and taking a shortest path to $e$. For any labeling $\ell(\theta)$, we define the active arcs $E^{\prime}_{\ell(\theta)}$ and the resetting arcs $E^{*}_{\ell(\theta)}$ as
[…]
@@
2.2 Equilibria
An equilibrium (also referred to as a dynamic equilibrium or a Nash flow over time) is a feasible flow over time in which almost all flow particles travel along a shortest path from $s$ to $t$, i.e., only along active arcs. In this case, resetting arcs are always active and the active and resetting arcs are characterized as follows (see [CCL15, Proposition 2]):
[…]
@@ 2.2 Equilibria
As proven in [CCL15], a feasible flow over time is an equilibrium if and only if $F^{+}_{e}(\ell_{v}(\theta))=F^{-}_{e}(\ell_{w}(\theta))$ for all arcs $e=vw$ and all departure times $\theta$. Define $x_{e}(\theta)\coloneqq F^{+}_{e}(\ell_{v}(\theta))$ for all $e$ and $\theta$. For an equilibrium flow, the derivative of $x_{e}$ at time $\theta$, which exists almost everywhere, can be interpreted as a flow describing what proportion of flow particles departing at time $\theta$ use arc $e$; $x^{\prime}$ is an $s$-$t$-flow of value $u_{0}$ almost everywhere. It has been shown [KS11] that for an equilibrium, the resulting pair $(\ell,x)$ satisfies the following for almost every $\theta$, setting $x^{\prime}=x^{\prime}(\theta)$, $\ell^{\prime}=\ell^{\prime}(\theta)$, $E^{\prime}=E^{\prime}_{\theta}$ and $E^{*}=E^{*}_{\theta}$.
[…]
@@
$\ell^{\prime}_{w}$ $\leq\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})\quad$ $\text{ for all }e=vw\in E^{\prime},$ (TF-3)
$\ell^{\prime}_{w}$ $=\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})$ $\text{ for all }e=vw\in E^{\prime}\text{ with }x^{\prime}_{e}>0,$ (TF-4)
@@
These are called the thin flow conditions (sometimes thin flows with resetting) for the configuration $(E^{\prime},E^{*})$. [CCL11] replaced (TF-3) with the stronger requirement
$\ell^{\prime}_{w}=\min_{e=vw\in E^{\prime}}\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})\qquad\text{ for all }w\in V\setminus\{s\},$ (TF-3)
and showed that the pair $(\ell,x)$ corresponding to a flow over time is an equilibrium if and only if these strengthened normalized thin flow conditions for configuration $(E^{\prime}_{\theta},E^{*}_{\theta})$ are satisfied by $(\ell^{\prime}(\theta),x^{\prime}(\theta))$ for almost every $\theta$. In particular, $\ell^{\prime}(\theta)$ and $x^{\prime}(\theta)$ exist for almost every $\theta$; we call such values of $\theta$ points of differentiability.
We will take the viewpoint throughout that an equilibrium is a pair $(\ell,x)$ whose derivative satisfies the normalized thin flow equations almost everywhere. Such a pair fully determines the flow over time $(f^{+},f^{-})$ of the equilibrium; for example, the queue volume is determined by $\tfrac{1}{\nu_{e}}z_{e}(\ell_{v}(\theta))=\max\{0,\ell_{w}(\theta)-\ell_{v}(\theta)-\tau_{e}\}$ for all $e=vw$ and $\theta\in\mathbb{R}_{\geq 0}$. Note that from this perspective, given $(\ell,x)$ satisfying the normalized thin flow conditions for all times $\theta\in[0,T]$, extending the equilibrium to later departure times only requires knowing $\ell(T)$; nothing about the earlier (with respect to departure time) behavior of the equilibrium is needed. Just $\ell$ alone captures all the truly important information; one could easily verify whether a given curve $\ell$ can be extended to an equilibrium, since given $\ell^{\prime}(\theta)$, it is easy to determine whether a matching $x^{\prime}(\theta)$ exists so that $(\ell^{\prime}(\theta),x^{\prime}(\theta))$ satisfy the normalized thin flow conditions.
The thin flow conditions, and the normalized thin flow conditions, are fully determined by the configuration $(E^{\prime},E^{*})$. However not all choices of $(E^{\prime},E^{*})$ are meaningful.
Definition 2.1.
We call a configuration $(E^{\prime},E^{*})$ with $E^{*}\subseteq E^{\prime}\subseteq E$ a valid configuration if
  1. (i)
    for every node $v\in V$ there is an $s$-$v$-path in $E^{\prime}$,
  2. (ii)
    every arc $e\in E^{*}$ lies on an $s$-$t$-path in $E^{\prime}$, and
  3. (iii)
    no arc of $E^{*}$ lies on a directed cycle in $(V,E^{\prime})$.
A vector $l\in\mathbb{R}^{V}$ is called feasible if $(E^{\prime}_{l},E^{*}_{l})$ is a valid configuration.
If $(\ell,x)$ is an equilibrium, then $\ell(\theta)$ is a feasible label for every $\theta$. Condition (i) is simply insisting that $\ell(\theta)$ does correspond to earliest arrival times from $s$, for some choice of nonnegative queue lengths on the arcs. Condition (ii) follows because by the definition of dynamic equilibrium, a flow particle at the back of a queue travels on an active $s$-$t$-path. The sum of transit times along a directed cycle is assumed to be strictly positive, meaning that $E^{\prime}_{\ell(\theta)}$ is acyclic and implying (iii).
[CCL15] showed that for any valid configuration $(E^{\prime},E^{*})$ with $E^{\prime}$ acyclic, the normalized thin flow conditions have a solution; from this they could deduce existence of equilibria. They also showed that the solution is label-unique: any other solution must have the same values for $\ell^{\prime}_{v}$ for all $v$. It follows that in an equilibrium, $\ell$ is a piecewise linear function; each linear segment is referred to as a phase of the equilibrium [KS11, CCL15].
We will require a slightly stronger result however: we will need existence and label-uniqueness for any valid configuration, without the restriction that $E^{\prime}$ be acyclic. (This is not strictly needed for our uniqueness and continuity result, but it will be important for our result on long-term behavior for general networks, as we will later see.) Existence was shown by Koch [Koc12, Theorem 6.61]. Label-uniqueness fails to hold, however, even with the inclusion of the normalization condition (TF-3). [^2] As an explicit example, consider a network $G$ with arcs $\{sv,vt,vw,wy,yz,zw,zs\}$ and capacities ${\color[rgb]{0,0,0}\nu}_{sv}=\tfrac{1}{2}$, ${\color[rgb]{0,0,0}\nu}_{e}=1$ for all other arcs $e$, with an inflow of $1$. It is easy to see that assigning $\ell^{\prime}_{s}=1$, $\ell^{\prime}_{v}=\ell^{\prime}_{t}=2$ and $\ell^{\prime}_{w}=\ell^{\prime}_{y}=\ell^{\prime}_{z}=\alpha$ is a solution to the normalized thin flow equations for configuration $(E,\emptyset)$ for all $\alpha\in[1,2]$.
Instead, we will use a slightly different strengthening of the thin flow conditions, and one that is completely equivalent if $E^{\prime}$ is acyclic. We will always consider a maximal solution to the thin flow equations (2), by which we mean a solution $(\ell^{\prime},x^{\prime})$ for which $\ell^{\prime}$ is pointwise maximal. Observe that this implies (TF-3): if $(\ell^{\prime},x^{\prime})$ is any solution to (2), and $w$ a node for which $\ell^{\prime}_{w}<\rho_{e}(\ell^{\prime}_{v},x^{\prime}_{e})$ for all $e=vw\in E^{\prime}$, then none of these arcs is flow carrying (by (TF-4)), and so increasing $\ell^{\prime}_{w}$ by a sufficiently small amount yields a new valid thin flow solution, showing that $(\ell^{\prime},x^{\prime})$ was not maximal. If $E^{\prime}$ is acyclic, the two notions are in fact identical, given that the normalized thin flow equations have a label-unique solution. But we now have label-uniqueness of maximal thin flow solutions for any valid configuration. The proof follows exactly the same lines as the proof in [CCL15] for the case where $E^{\prime}$ is acyclic, with minor modifications.
Lemma 2.2.
For any valid configuration $(E^{\prime},E^{*})$, maximal solutions to (2) are label-unique.
Proof.
Suppose for the sake of contradiction there are two maximal solutions to (2), namely $(\ell^{\prime},x^{\prime})$ and $(h^{\prime},y^{\prime})$ with $\ell^{\prime}\neq h^{\prime}$. Let $S\coloneqq\{v\in V\mid\ell_{v}^{\prime}>h^{\prime}_{v}\}$; this set is non-empty as otherwise $\ell^{\prime}$ would not be pointwise maximal.
Claim 2.3.
$x_{e}^{\prime}=y_{e}^{\prime}=0$ for all $e\in\delta^{+}(S)\cup\delta^{-}(S)$.
Proof.
Suppose first that $x_{e}^{\prime}>y_{e}^{\prime}\geq 0$ for some $e\in\delta^{+}(S)$. Then $\ell_{w}^{\prime}=\rho_{e}(\ell_{v}^{\prime},x_{e}^{\prime})>\rho_{e}(h_{v}^{\prime},y_{e}^{\prime})\geq h_{w}^{\prime}$, a contradiction as $w\not\in S$. Next, suppose $x_{e}^{\prime}<y_{e}^{\prime}$ and for some $e\in\delta^{-}(S)$. Then $h_{w}^{\prime}=\rho_{e}(h_{v}^{\prime},y_{e}^{\prime})\geq\rho_{e}(\ell_{v}^{\prime},x_{e}^{\prime})\geq l_{w}^{\prime}$ (using $y^{\prime}_{e}>0$ in the first equality), again a contradiction as $w\in S$.
The above yields that $x_{e}^{\prime}=y_{e}^{\prime}$ for all $e\in\delta^{+}(S)\cup\delta^{-}(S)$ as both $x^{\prime}$ and $y^{\prime}$ are static flows of the same flow value. Now, observe that $y_{e}^{\prime}>0$ for $e\in\delta^{-}(S)$ would still yield the contradiction $h^{\prime}_{w}\geq\ell^{\prime}_{w}$. Thus $x^{\prime}_{e}=y^{\prime}_{e}=0$ for all $e\in\delta^{-}(S)$, and as $s\not\in S$ this yields $x^{\prime}_{e}=y^{\prime}_{e}=0$ for all $e\in\delta^{+}(S)$ as well. ∎
Let
$k^{\prime}_{v}=\begin{cases}\ell_{v}^{\prime}\quad&\text{if }v\in S\\ h_{v}^{\prime}\quad&\text{otherwise,}\end{cases}\qquad\qquad\bar{x}^{\prime}_{e}=\begin{cases}x_{e}^{\prime}\quad&\text{if }e\in E[S]\\ y_{e}^{\prime}\quad&\text{otherwise.}\end{cases}$
We show that $(k^{\prime},\bar{x}^{\prime})$ is a thin flow for configuration $(E^{\prime},E^{*})$. By Section 2.2, $\bar{x}^{\prime}$ is an $s$-$t$-flow of value $u_{0}$ on $E^{\prime}$. (TF-2) and (TF-4) clearly remain valid, so we just have to make sure that (TF-3) is not violated by any edges crossing $S$. Thus, consider some entering edge $e=vw\in\delta^{-}(S)$. (TF-3) requires that $k_{w}^{\prime}\leq\rho_{e}(k_{v}^{\prime},0)=k_{v}^{\prime}$. This follows as $k_{w}^{\prime}=\ell_{w}^{\prime}\leq\rho(\ell_{v}^{\prime},0)=\ell_{v}^{\prime}\leq h_{v}^{\prime}=k_{v}^{\prime}$. Similarly, if we consider some leaving edge $e=vw\in\delta^{+}(S)$, (TF-3) requires that $k_{w}^{\prime}\leq\rho_{e}(k_{v}^{\prime},0)=k_{v}^{\prime}$. This follows as $k_{w}^{\prime}=h_{w}^{\prime}\leq\rho{\color[rgb]{0,0,0}{}_{e}}(h_{v}^{\prime},0)=h_{v}^{\prime}<l_{v}^{\prime}=k_{v}^{\prime}$. Thus, $(k^{\prime},\bar{x}^{\prime})$ is indeed a thin flow. Since $k^{\prime}\geq\ell^{\prime}$ and $k^{\prime}\geq h^{\prime}$, we obtain a contradiction to the assumption that $(\ell^{\prime},x^{\prime})$ and $(h^{\prime},y^{\prime})$ are distinct maximal solutions. ∎
From now on, we only consider maximal solutions to the thin flow equations. If $(\ell^{\prime},x^{\prime})$ is a maximal solution, we will refer to $\ell^{\prime}$, which is unique, as the thin flow direction.
The following simple fact will be useful.
Lemma 2.4.
For any valid configuration $(E^{\prime},E^{*})$ and corresponding maximal thin flow solution $(\ell^{\prime},x^{\prime})$, $\ell^{\prime}_{v}>0$ for all $v\in V$. Hence $x^{\prime}_{e}>0$ for every $e\in E^{*}$.
Proof.